### v2

# Fast-DDPM: Fast Denoising Diffusion Probabilistic Models for Medical Image-to-Image Generation

Abstract
Denoising diffusion probabilistic models (DDPMs) have achieved unprecedented success in computer vision. However, they remain underutilized in medical imaging, a field crucial for disease diagnosis and treatment planning. This is primarily due to the high computational cost associated with (1) the use of large number of time steps (e.g., 1,000) in diffusion processes and (2) the increased dimensionality of medical images, which are often 3D or 4D. Training a diffusion model on medical images typically takes days to weeks, while sampling each image volume takes minutes to hours. To address this challenge, we introduce Fast-DDPM, a simple yet effective approach capable of improving training speed, sampling speed, and generation quality simultaneously. Unlike DDPM, which trains the image denoiser across 1,000 time steps, Fast-DDPM trains and samples using only 10 time steps. The key to our method lies in aligning the training and sampling procedures to optimize time-step utilization. Specifically, we introduced two efficient noise schedulers with 10 time steps: one with uniform time step sampling and another with non-uniform sampling. We evaluated Fast-DDPM across three medical image-to-image generation tasks: multi-image super-resolution, image denoising, and image-to-image translation. Fast-DDPM outperformed DDPM and current state-of-the-art methods based on convolutional networks and generative adversarial networks in all tasks. Additionally, Fast-DDPM reduced the training time to 0.2$\times$ and the sampling time to 0.01$\times$ compared to DDPM. Our code is publicly available at: https://github.com/mirthAI/Fast-DDPM.
1 Introduction
Diffusion models [9, 18, 31, 32] have become powerful tools for high-quality image generation [5]. The forward diffusion process incrementally adds noise to a high-quality image until it becomes pure Gaussian noise. An image denoiser is then trained to learn the reverse process, progressively removing noise from random Gaussian noise until it reconstructs a noise-free image. This enables the generation of new high-quality images that match the distribution of the training dataset. Beyond computer vision, diffusion models are increasingly applied to medical imaging [35, 20, 29, 11, 15] for improved disease diagnosis, treatment planning, and patient monitoring.
Training and sampling diffusion models are computationally expensive and time-consuming, especially for medical images. This is due to (1) the large number of discrete time steps needed to approximate the continuous diffusion process, (2) the need for large training datasets to model complex image distributions, and (3) the increased dimensionality of medical images from 2D to 3D and 4D. For 2D images, training can take days on a single GPU, while sampling can take several minutes per image [31]. These times scale significantly with 3D and 4D medical imaging volumes, posing challenges to practical usability. This extensive training time limits experimentation with different model architectures and hyperparameters, while the sampling time of minutes to hours per image volume limits the real-time clinical applications and the hospital throughput.
Current research mainly focuses on accelerating the sampling process, using either training-free or training-based methods [41]. Training-free methods utilize efficient numerical solvers for stochastic or ordinary differential equations to reduce the number of time steps required for sampling. This approach can decrease the number of sampling time steps from 1,000 to 50-100, without compromising the quality of the generated images. Alternatively, training-based methods, such as progressive knowledge distillation [30, 26], can further reduce the sampling steps to as few as 10. However, this approach requires the training of additional student models, which considerably increases the overall training cost and may not be practical for medical image analysis.
In this paper, we propose Fast-DDPM, a model designed to reduce both training and sampling times by optimizing time-step utilization. We observed that the majority of time steps involved in training are not utilized by faster diffusion samplers during sampling. For instance, a Denoising Diffusion Probabilistic Model (DDPM) trained with 1,000 time steps is often sampled by the Denoising Diffusion Implicit Model (DDIM) using only 100 time steps. This practice is resource-inefficient, as the image denoiser is trained on 900 time points that are ultimately skipped during sampling. To address this inefficiency in training, we propose training and sampling the DDPM model at the same 10 time steps, as illustrated in Figure 1.
Figure 1: Fast-DDPM uses significantly fewer time steps than DDPM. For illustration purposes, we used only 3 time steps for Fast-DDPM in this example.
We evaluated Fast-DDPM on three medical image-to-image tasks, including multi-image super-resolution for prostate MRI, denoising low-dose CT scans, and image-to-image translation for brain MRI. Fast-DDPM achieved state-of-the-art performance across all three tasks. Notably, Fast-DDPM is approximately 100 times faster than the DDPM model during sampling and about 5 times faster during training.
2 Related Work
Fast Sampling of Diffusion Models.
Current methods to accelerate sampling of diffusion models can be classified as training-free and training-based. Training-free methods focus on developing diffusion solvers to efficiently solve reverse stochastic differential equations or their equivalent ordinary differential equations (ODEs). For instance, the DDIM sampler [31] models the reverse diffusion process using an ODE, significantly reducing the sampling time steps to 100 without impacting generation quality. DPM-Solver [22] introduces a high-order solver for diffusion ODEs based on an exact solution formulation that simplifies to an exponentially weighted integral, achieving high-quality sample generation with just 10-50 steps. A common training-based approach is knowledge distillation, where a teacher diffusion model is initially trained with a large number of time steps, followed by the sequential training of multiple student models with progressively fewer time steps, mirroring the teacher model’s behavior [30, 26]. This method reduces sampling time steps from 1,000 to 16 while maintaining sampling quality.
Diffusion Models in Medical Imaging.
Diffusion models have been applied to various medical imaging tasks, including segmentation [11, 16, 39], anomaly detection [37, 13, 40], denoising [7, 4, 24], reconstruction [21, 8, 1], registration [15], super-resolution [4, 24, 35, 38], and image-to-image translation [23, 29, 17]. Among these, the DDPM model is the most widely used in medical imaging due to its simplicity. Diffusion models can be categorized as 2D, 3D, or 4D based on input image dimensions. Currently, 2D models are the predominant choice in medical imaging due to their ease of implementation and lower memory requirements. However, when implementing diffusion models for 3D and 4D applications, training can take significantly longer, lasting anywhere from weeks to months. Improving the training and sampling processes is crucial for a broader adoption of diffusion models in medical imaging.
3 Methods
3.1 Background
Diffusion Models.
Starting from a noise-free image $x_{0}$, the forward diffusion process is a continuously-time stochastic process from time $t=0$ to $t=1$, incrementally introducing noise to the image $x_{0}$ until it becomes pure Gaussian noise at $t=1$. The distributions of the intermediate noisy images $x(t)$ are given by [18]:
$q(x(t)|x_{0})=\mathcal{N}(\alpha(t)x_{0},\sigma^{2}(t)\mathbb{I})$ (1)
where $x(0)=x_{0}$, $\alpha(t)$ and $\sigma(t)$ are differentiable functions, and the signal-to-noise ratio function $\text{SNR}(t):=\frac{\alpha^{2}(t)}{\sigma^{2}(t)}$ decreases monotonically from +$\infty$+ ∞ at $t=0$ to 00 at $t=1$. An equivalent formulation of $x(t)$ is given by the following stochastic differential equation (SDE):
$dx(t)=f(t)x(t)dt+g(t)dW(t)$ (2)
where $f(t)=\frac{\alpha^{\prime}(t)}{\alpha(t)}$, $g(t)=\sqrt{(\sigma^{2}(t))^{\prime}-2\frac{\alpha^{\prime}(t)}{\alpha(t)}\sigma^{2}(t)}$, and $W(t)$ is a standard Wiener process from $t=0$ to $t=1$. It has been shown that the reverse process of $x(t)$ is given by the following reverse-time SDE [32]:
$dx(t)=[f(t)-g^{2}(t)\nabla_{x(t)}\log p(x(t))dt+g(t)d\bar{W}(t)$ (3)
where $x(1)=\mathcal{N}(0,\mathbb{I})$, $d\bar{W}(t)$ the standard Wiener process backwards from $t=1$ to $t=0$, and $p(x(t))$ denote the probability density of $x(t)$. If the score function $\nabla_{x(t)}\text{log}p(x(t))$ is known for every $t\in[0,1]$, one can solve the reverse-time SDE to sample new images.
We can train a model $s_{\theta}(x(t),t)$ to estimate the score function using the following loss function:
$\underset{t\in[0,1],x_{0}\sim p_{\text{data}},x(t)\sim q(x(t)|x_{0})}{\mathbb{E}}\lambda(t)\left\|s_{\theta}(x(t),t)-\nabla_{x(t)}\log p(x(t)|x_{0})\right\|_{2}^{2}$ (4)
where $p(x(t)|x_{0})$ denote the density function of $x(t)$ generated in Eq. 1 and $\lambda(t)$ is a scalar-valued weight function. The training of $s_{\theta}(x(t),t)$ is straightforward since $q(x(t)|x_{0})$ is a Gaussian distribution and its score function has a closed form:
$\nabla_{x(t)}\text{log}p(x(t)|x_{0})=-\frac{x(t)-\alpha(t)x_{0}}{\sigma^{2}(t)}=-\frac{\epsilon}{\sigma(t)}$ (5)
where $\epsilon$ is a random noise sample from $\mathcal{N}(0,\mathbb{I})$ to generate $x(t)$ from $x_{0}$. Equivalently, we can use a U-Net denoiser $\epsilon_{\theta}(x(t),t)$ to estimate the random noise $\epsilon$ (i.e., the score function scaled by $-\sigma(t)$). Following [9], in this paper, we trained the U-Net denoiser using a simple loss function:
$\underset{i\in[{1,\cdots,T}],x_{0}\sim p_{\text{data}},\epsilon\sim\mathcal{N}(0,\mathbb{I})}{\mathbb{E}}[\epsilon-\epsilon_{\theta}(\alpha(t)x_{0}+\sigma(t)\epsilon,t)]^{2}$ (6)
where $T$ is the total number of time steps and $t=\frac{i}{T}$.
Linear-$\beta$ Noise Scheduler.
Any noise scheduler, defined by $\alpha(t)$ and $\sigma(t)$), should ensure that $\text{SNR}(1)=\frac{\alpha^{2}(1)}{\sigma^{2}(1)}\approx 0$ so that $x(1)$ is pure Gaussian noise. The variance-preserving noise schedulers [18] meet this requirement, where $\alpha(t)$ is a monotonically decreasing function, $\alpha(0)=1$, $\alpha(1)=0$, and $\alpha^{2}(t)+\sigma^{2}(t)=1$ for all $t$. This formulation guarantees that the variance of any latent image $x(t)$ is bounded. The most popular variance-persevering noise scheduler is the linear-$\beta$ noise scheduler used in the DDPM model [9]. In this approach, the forward diffusion process was defined as a Markovian process at discrete time steps $i\in\{1,2,\cdots,T\}$ using the following transition kernel:
$q\left(x\left(\frac{i}{T}\right)\middle|x\left(\frac{i-1}{T}\right)\right)=\mathcal{N}\left(\sqrt{1-\beta\left(\frac{i}{T}\right)}x\left(\frac{i-1}{T}\right),\beta\left(\frac{i}{T}\right)\mathbb{I}\right)$ (7)
where $t=\frac{i}{T}$, $\beta(t)=0.0001+(0.02-0.0001)t$, and the corresponding $\alpha$ noise scheduler $\alpha(t)$ at discrete time steps is given by
$\alpha\left(\frac{i}{T}\right)=\prod_{j=1}^{i}\left(1-\beta\left(\frac{j}{T}\right)\right)$ (8)
To establish feasible noise schedulers using Eq. 8, the total number of time steps $T$ should be carefully chosen. Figure 2 plots $\alpha^{2}(t)$ for different choices of $T$:
Figure 2: Impact of the number of time steps, $T$, on the noise scheduler.
100, 500, 1000, 2000, and 10000. As we can see, the total number of time steps cannot be too small (e.g., $T=100$), as this results in the endpoint of the forward diffusion process $x(1)$ not being purely Gaussian. Conversely, if the total number of time steps is too large (e.g., $T=10000$), $x(t)$ becomes pure Gaussian noise at an early time point (e.g., $t=0.2$). This makes the training of $\epsilon_{\theta}(x(t),t)$ at time steps $t>0.2$ unnecessary since there is no change in the noise level of $x(t)$ beyond this point. The plots in Figure 2 suggest that a good choice of $T$ lies between 500 and 2000. This explains why a linear-$\beta$ noise scheduler with $T=1000$ is currently the most popular choice [9]. In other words, we cannot reduce the number of time steps to 10 using Eq. 8. In this study, we first define the continuous function $\alpha^{2}(t)$ and then sample it at discrete time points.
3.2 Fast Denoising Diffusion Probabilistic Model (Fast-DDPM)
We have developed the Fast-DDPM model to accelerate the training and sampling of DDPM by utilizing task-specific noise schedulers with only 5-10 time steps.
Task-Specific $\alpha$ Noise Scheduler.
Previous studies indicate that images of different sizes require different noise schedulers for optimal results [3, 10]. Inspired by this, we propose two $\alpha$ noise schedulers for various medical image-to-image tasks. We first define a smooth, monotonically decreasing function $\alpha^{2}(t)$ from $t=0$ to $t=1$, with boundary conditions $\alpha^{2}(0)=1$ and $\alpha^{2}(1)=0$. In this paper, we use $\alpha^{2}(t)$ as defined by the linear-$\beta$ noise scheduler with $T=1,000$ in Eq.8 (see the green curve in Figure 2). Unlike the linear-$\beta$ noise scheduler used in the DDPM model, which uniformly samples 1,000 time points between [0,1], our first proposed noise scheduler uniformly samples 10 time steps between [0,1] (Figure 3(a)). The second proposed noise scheduler non-uniformly samples 10 time points between [0,1], with denser sampling at higher noise levels (Figure 3(b)). This formulation of noise schedulers satisfies all the desirable properties outlined earlier. Due to the reduced training and sampling time, users have the opportunity to try both noise schedulers for each image-to-image task to determine which one fits their task better.
(a) The Uniform Sampling
(b) The Non-Uniform Sampling
Figure 3: Proposed $\alpha$ noise schedulers with uniform and non-uniform sampling between [0,1].
We provide a rationale for our design. Previous works [32, 31] have demonstrated that the sampling process is equivalent to solving a reverse-time ODE of the image flow starting from $x(1)\sim\mathcal{N}(0,\mathbb{I})$. In a related context of diffeomorphic image registration, 10 time points are sufficient to solve an ODE for accurately estimating a complex time-dependent image flow between two images [28, shao]. Since we can sample at only 10 discrete time points during inference, we hypothesize that we only need to train the denoiser $\epsilon_{\theta}$ at those 10 time steps, not necessarily at all the 1,000 time points.
Training and Sampling of Fast-DDPM
In this paper, Fast-DDPM is applied to conditional image-to-image generation. For each image $x_{0}$ sampled from $p_{\text{data}}(x_{0})$, we assume there are one or more associated condition images, denoted by $c$. Here, $c$ may represent either a single image or multiple images, depending on the specific application. We denote the joint distribution of $x_{0}$ and $c$ as $p_{\text{joint}}(x_{0},c)$ and the marginal distribution of the conditional image as $p_{c}(c)$. We will train a conditional image denoiser $\epsilon_{\theta}(x(t),c,t)$ that takes the condition $c$ as an additional input to guide the estimation of the score function at each time $t$. We used the following simple loss function that assigns equal weights to different times $t$:
$\underset{i,(x_{0},c),\epsilon}{\mathbb{E}}||\epsilon-\epsilon_{\theta}(\alpha(t)x_{0}+\sigma(t)\epsilon,c,t)||^{2}$ (9)
where $T=10$, $i\in\{1,2,\cdots,T\}$, $t=\frac{i}{T}$, $(x_{0},c)\sim p_{\text{joint}}(x_{0},c)$, and $\epsilon\sim\mathcal{N}(0,\mathbb{I})$.
The sampling process starts by randomly sampling a Gaussian noise $x(1)\sim\mathcal{N}(0,\mathbb{I})$ and a condition image $c\sim p_{c}(c)$. The DDIM sampler [31] was used to solve the reverse-time ODE to obtain the noise-free output $x_{0}$ that corresponds to $c$. We summarize the training and sampling processes of Fast-DDPM in Algorithm 1 and Algorithm 2, respectively.
1 repeat
2       $i\in\{1,...,T\}$
3       $(x_{0},c)\sim p_{\text{joint}(x_{0},c)}$
4       $\epsilon\sim\mathcal{N}(0,\mathbb{I})$
5       $t=\frac{i}{T}$
6       Gradient descent: $\nabla_{\theta}||\epsilon-\epsilon_{\theta}(\alpha(t)x_{0}+\sigma(t)\epsilon,c,t)||^{2}$
7until converged;
Algorithm 1 Training Fast-DDPM
1 $x(1)\sim\mathcal{N}(0,\mathbb{I})$
2 $c\sim p_{\text{c}}(c)$
3 for $i=T,\cdots,1$ do
4       $t=\frac{i}{T}$
5       $x(t-\frac{1}{T})=\frac{\alpha(t-\frac{1}{T})}{\alpha(t)}x(t)+[\sigma(t-\frac{1}{T})-\frac{\alpha(t-\frac{1}{T})}{\alpha(t)}\sigma(t)]\epsilon_{\theta}\big{(}x(t),c,t)\big{)}$
6 end for
return $x(0)$
Algorithm 2 Sampling Fast-DDPM
4 Experiments
We evaluated our Fast-DDPM model on three medical image-to-image tasks: multi-image super-resolution, image denoising, and image-to-image translation.
4.1 Datasets
Prostate MRI Dataset for Multi-Image Super-Resolution.
We used T2-weighted (T2w) prostate MRI scans from the publicly available Prostate-MRI-US-Biopsy dataset [33]. The in-plane resolution of each MRI is 0.547mm $\times$ 0.547mm, with a 1.5mm distance between adjacent slices. An image triplet, comprising three consecutive MRI images, serves as a single training or testing example. In this setup, the first and third slices serve as inputs to the model, while the middle slice is considered as the ground truth. We used a total of 6979 image triplets from 120 MRI volumes for training and 580 image triplets from 10 MRI volumes for testing. All MRI slices were resized to $256\times 256$ and normalized to the range [-1, 1].
Low-Dose and Full-Dose Lung CT Dataset for Image Denoising.
We used paired low-dose and normal-dose chest CT image volumes from the publicly available LDCT-and-Projection-data dataset [25]. The normal-dose CT scans were acquired at routine dose levels, while the low-dose CT scans were reconstructed using simulated lower dose levels, specifically at 10% of the routine dose. We randomly selected 38 patients for training and the 10 patients for evaluation, comprising 13,211 and 3,501 2D low-dose and normal-dose CT image pairs, respectively. All images were resized to 256 $\times$ 256 and normalized to [-1, 1].
BraTS Brain MRI Dataset for Image-to-Image Translation.
We used registered T1-weighted (T1w) and T2-weighted (T2w) MR images from the publicly available BraTS 2018 dataset [27]. All images were resampled to a resolution of 1 mm³ and skull-stripped. Our focus was on converting T1w MRI into T2w MRI. The dataset, obtained from [19], comprised 5760 pairs of T1w and T2w MR images for training and 768 pairs for testing. All MRI slices were padded from 224 $\times$ 224 to 256 $\times$ 256 and normalized to [-1, 1].
4.2 Evaluation Metrics
In all three image-to-image generation tasks, we used the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM) to evaluate the similarity between the generated image and the ground truth image. PSNR measures the ratio between the maximum possible value (power) of the ground truth image and the power of distorting noise:
$PSNR=20\cdot\log_{10}\left(\frac{MAX_{I}}{\sqrt{MSE}}\right)$ (10)
where $MAX_{I}$denotes the maximum possible pixel value of the image and MSE denotes the mean squared error. SSIM [36] is a widely used metric that measures the structural similarity between two images and aligns better with human perception of image quality. SSIM is defined as:
$SSIM(x,\hat{x})=\frac{2\mu_{x}\mu_{\hat{x}}+c_{1}}{\mu^{2}{x}+\mu^{2}{\hat{x}}+c_{1}}\cdot\frac{2\sigma_{x\hat{x}}+c_{2}}{\sigma^{2}{x}+\sigma^{2}{\hat{x}}+c_{2}}$ (11)
where $\mu_{x},\mu_{\hat{x}}$ and $\sigma_{x},\sigma_{\hat{x}}$ are the means and standard deviations of image $x$ and image $\hat{x}$, respectively; $\sigma_{x\hat{x}}$ denotes the covariance of $x$ and $\hat{x}$. The value of $c_{1}$ is $(k_{1}L)^{2}$ and $c_{2}=(k_{2}L)^{2}$, where $k_{1}$ is 0.01, $k_{2}$ is 0.03, and $L$ denotes the largest pixel value of the image $x$.
4.3 Training Details
Fast-DDPM adopted the hyperparameter settings and the UNet architecture from DDIM [31]. For the image super-resolution task, we used the noise scheduler with non-uniform sampling, and for the denoising and translation tasks, we used the noise scheduler with uniform sampling. We used a training batch size of 16 and trained the Fast-DDPM model for 400,000 iterations and the DDPM model for 2 million iterations. The Adam optimizer with a learning rate of $2\times 10^{-4}$ was used for training. All experiments were conducted on a computing node with 4 NVIDIA A100 GPUs, each with 80GB memory. We used Python 3.10.6 and PyTorch 1.12.1 for all experiments.
5 Results
Given that DDPM is the most widely used diffusion model in medical imaging, we focused on comparing Fast-DDPM against DDPM and current state-of-the-art methods for each task.
5.1 Multi-Image Super-Resolution
Results in Table 1 show that Fast-DDPM outperformed DDPM, miSRCNN [6] (based on convolutional neural networks), and miSRGAN [34] (based on generative adversarial networks). Notably, Fast-DDPM significantly improved DDPM’s performance, increasing PSNR from 25.3 to 27.1 and SSIM from 0.83 to 0.89. Fast-DDPM also significantly improved computational efficiency compared to DDPM, reducing training time from 136 hours to 26 hours and per volume inference time from 3.7 minutes to 2.3 seconds.
Table 1: Comparison of the performance and inference time per image volume (an average of 58 slices) for various multi-image super-resolution methods.
Method PSNR SSIM Training time Inference Time
miSRCNN [6] 26.5 0.87 1 h 0.01s
miSRGAN [34] 26.8 0.88 40 h 0.04 s
DDPM [9] 25.3 0.83 136 h 3.7 m
Fast-DDPM 27.1 0.89 26 h 2.3 s
Figure 4 shows the results of various image super-resolution methods on a representative subject. The models received the previous and next slices as inputs, with the center slice serving as the ground truth. It is evident that the miSRCNN and miSRGAN methods failed to accurately reconstruct the shadowed area highlighted by the arrows. In contrast, Fast-DDPM most effectively reconstructed the information missing between the previous and next 2D slices. Interestingly, although DDPM is qualitatively inferior to miSRCNN and miSRGAN, it visually outperformed both methods.
Figure 4: Qualitative results of the MR multi-image super-resolution task.
5.2 Image Denoising
The results presented in Table 2 indicate that Fast-DDPM significantly outperformed the DDPM model, as well as two other prominent CT denoising methods, REDCNN [2] and DU-GAN [12]. In terms of computational efficiency, Fast-DDPM reduced the inference time by approximately 100-fold and the training time by 5-fold compared to DDPM. Notably, the performance of the original DDPM model was inferior to that of REDCNN and DU-GAN. By implementing an efficient training strategy, Fast-DDPM significantly improve the PSNR and SSIM, achieving state-of-the-art performance.
Table 2: Comparison of the performance and inference time per image volume (an average of 360 slices) for various CT image denoising methods.
Method PSNR SSIM Training time Inference Time
REDCNN [2] 36.4 0.91 3 h 0.5 s
DU-GAN [12] 36.3 0.90 20 h 3.8 s
DDPM [9] 35.4 0.87 141 h 21.4 m
Fast-DDPM 37.5 0.92 26 h 12.5 s
Figure 5 shows the low-dose, normal-dose, and predicted normal-dose CT images by various models of a representative subject. The zoomed-in region, highlighted by the red rectangular boxes, demonstrate the capability of our Fast-DDPM model to preserve detailed structures in the denoised images compared to other methods. Notably, the lung fissure appears clear and sharp in the image generated by Fast-DDPM, while the other methods exhibit noticeable blurring and loss of details in this region. A pulmonary fissure is a boundary between the lobes in the lungs. The preserved details of the fissure allow for automated segmentation of the lung fissure, facilitating further analysis of lung disease at a lobar level.
Figure 5: Qualitative results of the CT image denoising task.
5.3 Image-to-Image Translation
Results in Table 3 show that Fast-DDPM and DDPM outperformed leading methods in the image-to-image translation task, including Pix2Pix [14] and RegGAN [19]. Although DDPM achieved performance comparable to Fast-DDPM, Fast-DDPM significantly reduced the training time from 135 hours to 27 hours and the sampling time from 22.2 minutes per batch to 13.2 seconds per batch. This highlights the practical advantages of Fast-DDPM in real-world applications. Moreover, extending the training period for Fast-DDPM (denoted as Fast-DDPM*) enhanced its SSIM score from 0.89 to 0.91.
Table 3: Comparison of the performance and averaged inference time per batch (360 slices) for various image-to-image translation methods.
Method PSNR SSIM Training time Inference Time
Pix2Pix [14] 25.6 0.85 6 h 3.3 s
RegGAN [19] 26.0 0.86 9 h 3.1 s
DDPM [9] 26.3 0.89 135 h 22.2 m
Fast-DDPM 26.3 0.89 27 h 13.2 s
Fast-DDPM* 26.3 0.91 133 h 13.2 s
Figure 6: Qualitative results of the T1w MRI to T2w MRI translation task.
Figure 6 shows the T2w MR images generated by various image-to-image translation methods for a representative subject. Notably, predictions by the Pix2Pix and RegGAN models are blurry and fail to reconstruct the intricate structures in the brain region highlighted by the arrows. Although the DDPM model recovered some of the details in the MRI, its reconstructed T2 MRI has a relatively lower signal-to-noise ratio. The Fast-DDPM model demonstrates its ability to generate high-quality translated images with better structural details and less noise.
5.4 Ablation Study
Impact of the Number of Time Steps.
We investigated the impact of the number of time steps on the performance of Fast-DDPM. Specifically, we conducted experiments using 5, 10, 20, 50, 100, and 1000 time steps across all three datasets. The results in Table 4 indicate that the number of time steps significantly affects the performance of Fast-DDPM in different image-to-image tasks. Optimal performance for image super-resolution and image denoising is achieved with 10 time steps. For the image-to-image translation task, where the condition ($c$) and the output ($x_{0}$) have significantly different intensity levels, the best performance is achieved using 100 time steps. For all tasks, performance declines with more than 100 time steps. These findings underscore the importance of carefully tuning the number of time steps for specific tasks to balance computational efficiency and image quality. Future work could explore adaptive mechanisms for dynamically adjusting time steps based on input complexity and task requirements.
Table 4: The impact of the number of time steps on different image-to-image tasks.
Number of Time Steps Super-Resolution Denoising Translation
PSNR SSIM PSNR SSIM PSNR SSIM
5 26.9 0.88 37.4 0.92 25.8 0.91
10 27.1 0.89 37.5 0.92 26.3 0.91
20 26.7 0.88 37.4 0.91 26.5 0.91
50 26.2 0.87 36.9 0.91 26.5 0.92
100 25.8 0.86 36.5 0.90 26.6 0.92
1000 25.3 0.83 35.4 0.87 26.3 0.89
Impact of Noise Scheduler.
Table 5 illustrates the impact of the two proposed noise schedulers on the performance of Fast-DDPM. The uniform noise scheduler is a better choice for image denoising and image-to-image translation tasks, while the non-uniform noise scheduler performed better in the super-resolution task. These findings highlight the importance of selecting an appropriate noise scheduler tailored to the specific image-to-image task. This adaptability enhances Fast-DDPM’s performance across various medical image processing applications, demonstrating its versatility and effectiveness in handling diverse imaging challenges.
Table 5: The impact of two proposed noise schedulers on different image-to-image tasks.
Noise Scheduler Super-Resolution Denoising Translation
PSNR SSIM PSNR SSIM PSNR SSIM
Uniform 26.6 0.88 37.5 0.92 26.3 0.91
Non-uniform 27.1 0.89 37.3 0.91 26.1 0.90
6 Discussion and Conclusion
This paper introduces Fast-DDPM, a simple yet effective approach that accelerates both the training and sampling of DDPM while significantly improving its performance. Our evaluation demonstrates that Fast-DDPM achieves state-of-the-art performance across various medical image-to-image generation tasks.
Limitations and Future Work
The noise schedulers used in this study were manually designed and optimized for the specific tasks, this approach may not be optimal for all types of medical imaging tasks. Future work should focus on developing adaptive noise schedulers that can dynamically adjust based on the complexity of the input data and the specific requirements of the imaging task.
Broader Impact
The advancements introduced by Fast-DDPM have significant implications for the field of medical imaging. The dramatic reduction in computational resources required for training and sampling makes diffusion models more accessible and practical for real-world clinical applications. This can lead to faster and more accurate diagnosis and treatment planning, ultimately improving patient outcomes.